OpenStudy (anonymous):

Find the quotient of 21 x squared y minus 15 x y squared plus 9 x y all over 3 x y

OpenStudy (jackellyn):

For 12x^2y/ 3xy 21/3 =7 x^2 -x^1 =x y-y =0 Which gives: 7x For -15xy^2/3xy -15/3 = -5 x-x=0 y^2-y =1 Which gives: -5y For 9xy/3xy 9/3=3 xy-xy=0 Which gives: 3 Answer: 7x-5y +3